If the money supply is growing at 7 percent, the real interest rate is 3 percent, and the real of growth rate of GDP is 2 percent, what should the nominal interest rate be?
1
Expert's answer
2012-11-27T08:33:16-0500
ΔY=+2%
ΔM=+7%
ΔV=0%
r=3%

M/P=Y/V
(1+ΔM%)/ (1+ΔP%) = (1+ΔY%) / (1+ΔV%)
(1+7%)/ (1+ΔP%) = (1+2%) / (1+0%)
1.07/(1+ΔP%)=1.02
(1+ΔP%)=1.07/1.02≈1.049
π%=ΔP%≈4.9%